# DiTASK: Multi-Task Fine-Tuning with Diffeomorphic Transformations

### v2

@@ Abstract
Figure 1: Performance-efficiency trade-off in multi-task learning (MTL) across parameter-efficient fine-tuning methods on PASCAL MTL tasks. Our method DiTASK achieves superior relative improvement over single-task fine-tuning while using significantly fewer parameters compared to fixed-rank approaches (LoRA [19]) and other PEFT methods (Adapter [18]). The gray region indicates negative relative improvement versus single-task fine-tuning, illustrating the fundamental trade-off in MTL: methods with fewer parameters struggle to match single-task performance while achieving positive improvements typically requires substantial parameter overhead.
@@ 1 Introduction
Vision Transformers (ViTs) have emerged as compelling models for computer vision, achieving remarkable performance across various tasks by leveraging self-attention mechanisms [46, 12] and large-scale pre-training [40, 37] on extensive datasets like ImageNet-21k [10]. These models encode essential visual knowledge in their weight matrices through singular vectors that characterize the principal directions in the input space and output spaces [34, 38]. However, existing fine-tuning methods based on low-rank adaptation (LoRA) [19, 1] may not preserve these learned singular vectors. In Figure 2, we illustrate the importance of preserving the singular vectors; recovering a noisy image by adjusting only its singular values while keeping singular vectors yields superior performance than using LoRA, which may not preserve the singular vectors of the image. Singular vectors encode crucial image details and low-rank denoising methods can lose these details, resulting in poor denoising quality. Preserving singular vectors serves as an implicit regularizer. This motivating example suggests that useful adaptations can be achieved through singular value modulation and inspires us to design a mathematically-ground method to achieve this goal.
Figure 2: Image recovery with LoRA vs DiTASK. The preservation of singular vectors in DiTASK offers an improved recovery with a higher (better) peak signal-to-noise ratio (PSNR).
@@
(a) Input feature ${\mathbf{x}}$ is transformed through the weight matrix ${\mathbf{W}}_{A}$ obtained after modulating the singular values of pre-trained weight ${\mathbf{W}}$ using a neural diffeomorphism $f^{{\bm{\theta}}}$ to obtain updated features ${\mathbf{W}}_{A}{\mathbf{x}}$.
(b) We use two sets of DiTASK modules, a joint transformation module to learn task synergies, and a set of task-wise modules to adapt to each task. The outputs are used as multi-task inputs to the decoder.
Figure 3: Overview of our DiTASK within Multi-Task Learning (MTL). DiTASK leverages learned neural diffeomorphic transformations for both joint and task-wise adaptations. This approach achieves multi-task adaptation, preserving the singular vectors of pre-trained weights while enabling flexible, parameter-efficient fine-tuning.
@@ 2 Related Work
Our work lies at the intersection of four research areas: multi-task vision learning, parameter-efficient fine-tuning, neural network adaptation, and diffeomorphisms. We discuss each area with a focus on how they relate to weight matrix and singular value adaptation.
Multi-Task Learning in Vision.
Multi-task learning models in vision typically adopt an encoder-decoder structure, where a shared encoder captures common features, and task-specific decoders handle individual objectives [33]. This approach has proven particularly effective for dense prediction tasks, with different decoders handling semantic segmentation, depth estimation, and surface normal prediction while sharing a common feature backbone [24]. While recent work has extended this framework to Vision Transformers (ViTs) [12], these adaptations face unique challenges due to their large parameter count and the complex interactions between task-specific updates. To address these issues, DiTASK introduces an encoder adaptation strategy that modulates singular values while retaining pre-trained singular vectors, rather than modifying full weight matrices or adding task-specific layers. This leverages the encoder-decoder framework for efficient adaptation of the shared encoder.
Parameter-Efficient Fine-tuning Methods.
These methods reduce computational overhead for model adaptation using various strategies. Adapter-based approaches [18] insert small trainable modules between transformer layers. Visual Prompt Tuning (VPT) [22] introduces trainable parameters in ViT input tokens, distinguishing it from standard adaptation approaches. Compactor [31] and BitFit [56] add minimal trainable parameters, enabling lightweight gradients via Kronecker products of rank-one and low-rank matrices and fine-tuning only biases, respectively. Hyperformers [31] use hypernetworks to generate adapter layers for multi-task language learning, while Polyhistor [27] adapts this for dense vision tasks. VL-Adapter [42] introduces adapters for joint Vision-Language (VL) multi-task fine-tuning.
@@
Multi-Task Learning Dynamics and Challenges.
A recent study of Transformer training dynamics reveals an incremental learning phenomenon where weight update ranks gradually increase during training [5]. This finding challenges the effectiveness of fixed-rank adaptation methods in MTL settings. While MTLoRA [1] attempts to address this through task-agnostic and task-specific modules, it still operates within fixed low-rank constraints. Similarly, gradient interference mitigation techniques like PCGrad address conflicts between task updates but do not provide the flexibility needed for diverse task adaptations. In contrast, DiTASK aligns with natural learning dynamics by preserving the full-rank structure through singular vector retention while enabling task-specific adaptations via neural diffeomorphisms [14, 15], uniquely combining parameter efficiency with the adaptability needed for effective multi-task learning.
Diffeomorphisms in Neural Networks.
A bijective mapping $f:{\mathcal{M}}\to{\mathcal{N}}$ between two differentiable manifolds ${\mathcal{M}}$ and ${\mathcal{N}}$ is a diffeomorphism if both $f$ and its inverse $f^{-1}:{\mathcal{N}}\to{\mathcal{M}}$ are differentiable. Learning diffeomorphisms poses computational challenges, as early methods were constrained by complex, infinite-dimensional spaces [41], and later Markov Chain Monte Carlo methods remained demanding [2, 3, 58]. Freifeld et al. [14, 15] addressed these issues with Continuous Piecewise-Affine Based (CPAB) transformations, a finite-dimensional approach suited for precise 1D diffeomorphisms, ideal for neural activation functions. CPAB offers linear complexity, supports parallelization, and achieves sub-linear performance in practice [15]. Initially developed for alignment and regression, CPAB is now widely applied in neural networks. For instance, Detlefsen et al. [11] enhanced spatial transformer layers with CPAB, Martinez et al. [32] applied CPAB to temporal alignment, Weber and Freifeld [50] developed a CPAB-free loss for time-series analysis, Wang et al. [47] used CPAB for image animation, and Chelly et al. [8] proposed a trainable CPAB-based activation function.
@@ 3.1 Learned Weights in ViTs
Figure 4: DiTASK in $i$-th Swin Transformer Stage for Multi-Task Learning. Task-agnostic modules are applied in all blocks except the last, enabling shared feature learning across tasks. In the last block, Task-Specific modules capture task-dependent features.
ViTs encode visual knowledge in their weight matrices, which can be analyzed through Singular Value Decomposition (SVD). Assume we have a weight matrix that transforms inputs from $c_{1}$ to $c_{2}$ channels, denoted by $\mathbf{W}\in\mathbb{R}^{c_{2}\times c_{1}}$. Its SVD reads:
@@ 3.2 CPAB Diffeomorphisms for Weight Transformation
As we show later in Section 4, our DiTASK modifies singular values $\mathbf{\Sigma}$ while preserving the structure of weight matrices. To achieve that, transformations that are smooth, invertible, and monotone, are required. Continuous Piecewise Affine-Based (CPAB) [14, 15] diffeomorphisms satisfy these conditions, while being efficient in terms of parameters and computing diffeomorphisms. We now turn to define CPAB formally.
@@ 4.2 Adaptation with DiTASK
During fine-tuning, we freeze ${\mathbf{W}}$ and learn only the parameters ${\bm{\theta}}$ of the CPA velocity field. This allows DiTASK to adapt to multiple tasks using ${\mathcal{O}}({\mathcal{N}}_{\mathcal{P}})$ additional parameters (where typically ${\mathcal{N}}_{\mathcal{P}}\leq 128$), which is significantly smaller than methods with low-rank $r$ using ${\mathcal{O}}(r(c_{1}+c_{2}))$. We describe this process in the Figure 3(a), and compare the computational cost of DiTASK with other fine-tuning methods in the supplementary material.
Multi-Task Adaptation with DiTASK. To enable both multi and task-specific adaptation, we learn two sets of neural diffeomorphic transformations in every pre-trained weight matrix ${\mathbf{W}}$, shown in Figure 3(b), that include: (i) Joint Adaptation parameters, denoted by ${\bm{\theta}}_{j}$, to allow task synergies which are crucial for MTL [21, 45]; and (ii) Task-wise Adaptation parameters, denoted by $\{{\bm{\theta}}_{k}\}_{k=1}^{K}$ to enable learning nuances for each task.
@@ 5 Experiments
Table 1: Performance Comparison on PASCAL Context for Multi-Task Learning: This table summarizes the performance of our proposed DiTASK method compared to various fine-tuning approaches, including full fine-tuning, adapter-based methods, and fixed-rank approaches. Results are reported for four dense prediction tasks: Semantic Segmentation (SemSeg), Human Part Segmentation, Saliency Detection, and Surface Normal Estimation. Metrics include mIoU for segmentation tasks and rmse for normal estimation, with $\Delta m$ indicating the percentage improvement over single-task fine-tuning. Our method achieves superior performance with minimal trainable parameters, enabling efficient single inference across all tasks.
@@ 5.2 Results and Discussion
Our main results, summarized in Table 1, address (RQ1)-(RQ3) with respect to the effectiveness of DiTASK in MTL, as discussed below.
Bridging the Efficiency-Performance Gap. DiTASK achieves a $\mathbf{+3.22\%}$ relative improvement over single-task baselines using only 1.61M trainable encoder parameters. This significantly outperforms both decoder-only tuning ($-9.95\%$) and full fine-tuning ($+2.23\%$), demonstrating that selective modulation of singular values while preserving pre-trained feature directions effectively balances parameter efficiency and performance.
Comparison with Fixed Low-Rank Methods. By maintaining full-rank structure through singular value modulation, DiTASK achieves a $26.27\%$ improvement in $\Delta m$ over the state-of-the-art MTLoRA, using $\mathbf{4\times}$ fewer parameters (see Figure 5). Most fixed low-rank methods show negative $\Delta m$, as restricting updates to shared low-rank subspaces causes task interference. Allowing singular values to be modulated independently for each task, our approach mitigates this interference effectively. Similar conclusions can drawn for NYUD based on Table 4.
Parameter Efficiency and Task Performance. Among all evaluated parameter-efficient methods, DiTASK consistently yields positive improvements over single-task baselines, confirming that preserving pre-trained singular vectors with task-specific singular value modulation balances parameter efficiency and performance effectively.
Figure 5: Comparison of adaptation parameter budget of the shared encoder (excluding attention mask and layerNorm)
@@ 5.3 Ablation Study
Figure 6: Effect of task-specific and task-agnostic components
@@
Table 3: Computational efficiency comparison between DiTASK and MTLoRA. DiTASK achieves faster inference ($\approx 22\%$ reduction in batch runtime) and lower memory footprint while requiring 500×500$\times$500 × fewer adaptation parameters.
Method Batch Runtime Max GPU Mem # Adaptation (seconds) (Megabytes) Params MTLoRA 35.20$\pm$1.05 23509 4.86m DiTASK (ours) 27.31$\pm$5.29 22906 8.5k
Qualitative Comparison. In Figure 7 we present a visual comparison between MTLoRA and our DiTASK on the semantic segmentation task from PASCAL MTL. DiTASK demonstrates tighter and clearer boundaries, more consistent segmentation, and improved class separation. For example, in the dimly lit sofa scene in the second column, MTLoRA struggles to correctly predict the sofa boundary, while DiTASK offers a better prediction.
Method SemSeg Depth ${\Delta m(\%)}$ Trainable Params $(\textsc{mIoU})\uparrow$ $(\textsc{rmse})\downarrow$ (in M) Single Task 33.18 0.667 0 112.62 MTL - Dec. Only 28.37 0.832 -19.61 1.00 MTL - Full 35.29 0.734 -1.84 28.5 MTLoRA 37.18 0.635 +8.42 6.26 Single Task - DiTASK 44.01 0.644 +18.04 1.61 DiTASK (ours) 43.85 0.606 +20.65 1.61
Table 4: Comparison of DiTASK with MTLoRA on NYUD [39] MTL dataset for Semantic Segmentation (SemSeg) and Depth Estimation tasks. We use mIoU for segmentation and rmse for depth estimation as metrics.
@@ 6 Conclusion
Figure 7: Semantic segmentation predictions on the PASCAL MTL dataset with MTLoRA and our DiTASK
@@ Appendix B Gradient Analysis
In contrast, DiTASK leverages the singular value decomposition (SVD) of ${\mathbf{W}}={\mathbf{U}}\mathbf{\Sigma}{\mathbf{V}}^{\top}$, where $\mathbf{\Sigma}=\text{diag}(\sigma_{1},\cdots,\sigma_{p}),p=\min(c_{1},c_{2})$. DiTASK adapts ${\mathbf{W}}$ by learning transformations on the singular values $\mathbf{\Sigma}$ using a small set of parameters ${\bm{\theta}}$. This reduces the gradient storage requirement to ${\mathcal{N}}_{\mathcal{P}}+p+c_{1}c_{2}$, where ${\mathcal{N}}_{\mathcal{P}}$ is the number of intervals over which CPA velocity field is defined. Unlike LoRA, DiTASK avoids gradients tied to low-rank matrices, significantly reducing memory usage for tasks with high-rank requirements or large input dimensions.
@@ Appendix D Additional Ablations
VTAB Benchmark. To understand the single-task generalization capabilities of our method, we compare DiTASK and LoRA [19] using the ViT [12] backbone on the VTAB Benchmark [57]. Shown in Figure 9, DiTASK achieves competitive performance with an order of magnitude fewer learnable parameters presenting itself as a strong fine-tuning method for vision transformers.
@@
Additional Experiments. To understand the robustness of our design for different backbones, we evaluate DiTASK using the Pyramid Vision Transformer (PVT) [49] backbone against various parameter budgets of MTLoRA and LoRA. From Table 6, we conclude that our DiTASKachieves 2.5$\times$ improvement in multi-task performance over best performing baseline using 4.4$\times$ fewer trainable backbone parameters.

### v3

@@ 1 Introduction
Vision Transformers (ViTs) have emerged as compelling models for computer vision, achieving remarkable performance across various tasks by leveraging self-attention mechanisms [48, 12] and large-scale pre-training [42, 39] on extensive datasets like ImageNet-21k [10]. These models encode essential visual knowledge in their weight matrices through singular vectors that characterize the principal directions in the input space and output spaces [36, 40].
Figure 1: Performance–efficiency trade-off in multi-task learning (MTL) on PASCAL MTL tasks. DiTASK outperforms single-task fine-tuning with fewer parameters than fixed-rank methods (LoRA [19]) and other PEFT approaches (Adapter [18]). The gray region shows negative relative improvement over single-task tuning, highlighting the challenge of achieving gains under tight parameter budgets.
However, existing fine-tuning methods based on low-rank adaptation (LoRA) [19, 1] may not preserve these learned singular vectors. In Figure 2, we illustrate the importance of preserving the singular vectors; recovering a noisy image by adjusting only its singular values while keeping singular vectors yields superior performance than using LoRA, which may not preserve the singular vectors of the image. Singular vectors encode crucial image details, and low-rank denoising methods can lose these details, resulting in poor denoising quality. Preserving singular vectors serves as an implicit regularizer. This motivating example suggests that useful adaptations can be achieved through singular value modulation and inspires us to design a mathematically-ground method to achieve this goal.
Figure 2: Image recovery comparison of LoRA and DiTASK. Preserving singular vectors enables DiTASK to achieve a higher peak signal-to-noise ratio (PSNR).
@@
Figure 3: Overview of our DiTASK within Multi-Task Learning (MTL). (Left) Input features ${\mathbf{x}}$ are transformed via a modulated weight matrix ${\mathbf{W}}_{A}$, constructed by applying a neural diffeomorphism $f^{{\bm{\theta}}}$ to the singular values of a pre-trained weight matrix ${\mathbf{W}}$, resulting in updated features ${\mathbf{W}}_{A}{\mathbf{x}}$. (Right) We use two sets of DiTASK modules: a joint transformation module for learning task synergies, and task-specific modules for individual adaptations. The transformed outputs serve as inputs to a shared decoder. This setup enables flexible, parameter-efficient fine-tuning while preserving the singular vectors of pre-trained weights.
@@ 2 Related Work
Our work lies at the intersection of four research areas: multi-task vision learning, parameter-efficient fine-tuning, neural network adaptation, and diffeomorphisms. We discuss each area with a focus on how they relate to weight matrix transformations and singular value adaptation.
Multi-Task Learning in Vision. Multi-task learning models in vision typically adopt an encoder-decoder structure, where a shared encoder captures common features, and task-specific decoders handle individual objectives [35]. This approach has proven particularly effective for dense prediction tasks, with different decoders handling semantic segmentation, depth estimation, and surface normal prediction while sharing a common feature backbone [24]. While recent work has extended this framework to Vision Transformers (ViTs) [12], these adaptations face unique challenges due to their large parameter count and the complex interactions between task-specific updates. To address these issues, DiTASK introduces an encoder adaptation strategy that modulates singular values while retaining pre-trained singular vectors, rather than modifying full weight matrices or adding task-specific layers. This leverages the encoder-decoder framework for efficient adaptation of the shared encoder.
Parameter-Efficient Fine-tuning Methods. These methods reduce computational overhead for model adaptation using various strategies. Adapter-based approaches [18] insert small trainable modules between transformer layers. Visual Prompt Tuning (VPT) [22] introduces trainable parameters in ViT input tokens, distinguishing it from standard adaptation approaches. Compactor [31] and BitFit [58] add minimal trainable parameters, enabling lightweight gradients via Kronecker products of rank-one and low-rank matrices and fine-tuning only biases, respectively. Hyperformers [31] use hypernetworks to generate adapter layers for multi-task language learning, while Polyhistor [27] adapts this for dense vision tasks. VL-Adapter [44] introduces adapters for joint Vision-Language (VL) multi-task fine-tuning.
@@
Multi-Task Learning Dynamics and Challenges. A recent study of Transformer training dynamics reveals an incremental learning phenomenon where weight update ranks gradually increase during training [5]. This finding challenges the effectiveness of fixed-rank adaptation methods in MTL settings. While MTLoRA [1] attempts to address this through task-agnostic and task-specific modules, it still operates within fixed low-rank constraints. Similarly, gradient interference mitigation techniques like PCGrad address conflicts between task updates but do not provide the flexibility needed for diverse task adaptations. In contrast, DiTASK aligns with natural learning dynamics by preserving the full-rank structure through singular vector retention while enabling task-specific adaptations via neural diffeomorphisms [14, 15], uniquely combining parameter efficiency with the adaptability needed for effective multi-task learning.
Diffeomorphisms in Neural Networks. A bijective mapping $f:{\mathcal{M}}\to{\mathcal{N}}$ between two differentiable manifolds ${\mathcal{M}}$ and ${\mathcal{N}}$ is a diffeomorphism if both $f$ and its inverse $f^{-1}:{\mathcal{N}}\to{\mathcal{M}}$ are differentiable. Learning diffeomorphisms poses computational challenges, as early methods were constrained by complex, infinite-dimensional spaces [43], and later Markov Chain Monte Carlo methods remained demanding [2, 3, 60]. Freifeld et al. [14, 15] addressed these issues with Continuous Piecewise-Affine Based (CPAB) transformations, a finite-dimensional approach suited for precise 1D diffeomorphisms, ideal for neural activation functions. CPAB offers linear complexity, supports parallelization, and achieves sub-linear performance in practice [15]. Initially developed for alignment and regression, CPAB is now widely applied in neural networks. For instance, in spatial transformer layers with CPAB [11], temporal alignment [34], novel loss function for time-series analysis [52], image animation [49], geometry-guided fine-tuning [32] and trainable CPAB-based activation function [33, 8].
@@ 3.1 Learned Weights in ViTs
Figure 4: DiTASK in the $i$-th Swin Transformer stage for multi-task learning. Task-agnostic modules are applied in all blocks except the last, which uses task-specific modules to capture task-dependent features.
ViTs encode visual knowledge in their weight matrices, which can be analyzed through Singular Value Decomposition (SVD). Assume we have a weight matrix that transforms inputs from $c_{1}$ to $c_{2}$ channels, denoted by $\mathbf{W}\in\mathbb{R}^{c_{2}\times c_{1}}$, then its SVD reads:
@@ 3.2 CPAB Diffeomorphisms for Weight Transformation
As we show later in Section 4, our DiTASK modifies singular values $\mathbf{\Sigma}$ while preserving the structure of weight matrices. To achieve that, transformations that are smooth, invertible, and monotone, are required. Continuous Piecewise Affine-Based (CPAB) [14, 15] diffeomorphisms satisfy these conditions while being efficient in terms of parameters and computing diffeomorphisms. We now turn to define CPAB formally.
@@ 4.2 Adaptation with DiTASK
During fine-tuning, we freeze ${\mathbf{W}}$ and learn only the parameters ${\bm{\theta}}$ of the CPA velocity field. This allows DiTASK to adapt to multiple tasks using ${\mathcal{O}}({\mathcal{N}}_{\mathcal{P}})$ additional parameters (where typically ${\mathcal{N}}_{\mathcal{P}}\leq 128$), which is significantly smaller than methods with low-rank $r$ using ${\mathcal{O}}(r(c_{1}+c_{2}))$. We describe this process in the Figure 3, and compare the computational cost of DiTASK with other fine-tuning methods in the supplementary material.
Multi-Task Adaptation with DiTASK. To enable both multi and task-specific adaptation, we learn two sets of neural diffeomorphic transformations in every pre-trained weight matrix ${\mathbf{W}}$, shown in Figure 3, that include: (i) Joint Adaptation parameters, denoted by ${\bm{\theta}}_{j}$, to allow task synergies which are crucial for MTL [21, 47]; and (ii) Task-wise Adaptation parameters, denoted by $\{{\bm{\theta}}_{k}\}_{k=1}^{K}$ to enable learning nuances for each task.
@@ 5 Experiments
Table 1: Performance on PASCAL Context for Multi-Task Learning. Comparison of DiTASK with full fine-tuning, adapter-based, and fixed-rank methods across four dense prediction tasks: semantic segmentation, human part segmentation, saliency detection, and surface normal estimation. Metrics are mIoU ($\uparrow$) and rmse ($\downarrow$), with $\Delta m$ denoting relative improvement over single-task tuning. DiTASK achieves strong performance with minimal trainable parameters and supports joint inference across tasks.
@@ 5.2 Results and Discussion
Table 1 presents our main results, addressing (RQ1)–(RQ3) on the effectiveness of DiTASK for multi-task learning (MTL).
Bridging the Efficiency–Performance Gap. DiTASK achieves a $\mathbf{+3.22\%}$ relative improvement over single-task baselines using only 1.61M trainable encoder parameters. This outperforms both decoder-only tuning ($-9.95\%$) and full fine-tuning ($+2.23\%$), demonstrating that modulating singular values while preserving pre-trained feature directions enables an effective balance between parameter efficiency and performance.
Comparison with Fixed Low-Rank Methods. DiTASK outperforms the state-of-the-art MTLoRA by $26.27\%$ in $\Delta m$, using $\mathbf{4\times}$ fewer parameters (Figure 5). Most fixed-rank methods yield negative $\Delta m$ due to task interference from shared low-rank subspaces. By modulating singular values independently per task while preserving full-rank structure, DiTASK mitigates this interference. Similar trends are observed on NYUD (Table 4).
Parameter Efficiency and Task Performance. DiTASK is the only parameter-efficient method to consistently improve over single-task baselines, showing that modulating singular values while preserving pre-trained directions effectively balances efficiency and performance.
Figure 5: Comparison of adaptation parameter budget of the shared encoder, excluding LayerNorm parameters.
@@ 5.3 Ablation Study
Table 3: Comparison of computational efficiency between DiTASK and MTLoRA. DiTASK reduces batch runtime by $\approx 22\%$, lowers memory usage, and uses 500×500$\times$500 × fewer adaptation parameters.
Method Batch Runtime Max GPU Mem # Adaptation (seconds) (Megabytes) Params MTLoRA 35.20$\pm$1.05 23509 4.86m DiTASK 27.31$\pm$5.29 22906 8.5k
Qualitative Comparison. In Figure 6, we present a visual comparison between MTLoRA and our DiTASK on the semantic segmentation task from PASCAL MTL. DiTASK demonstrates tighter and clearer boundaries, more consistent segmentation, and improved class separation. For example, in the dimly lit sofa scene in the second column, MTLoRA struggles to correctly predict the sofa boundary, while DiTASK offers a better prediction.
[…]
Table 4: Comparison of DiTASK and MTLoRA on the NYUD [41] multi-task dataset for semantic segmentation (mIoU) and depth estimation (rmse).
@@ 6 Conclusion
Figure 6: Semantic segmentation predictions on PASCAL MTL using MTLoRA and DiTASK.
@@
Figure 7: Effect of task-specific and task-agnostic components
@@ Appendix B Gradient Analysis
In contrast, DiTASK leverages the singular value decomposition (SVD) of ${\mathbf{W}}={\mathbf{U}}\mathbf{\Sigma}{\mathbf{V}}^{\top}$, where $\mathbf{\Sigma}=\text{diag}(\sigma_{1},\cdots,\sigma_{p}),p=\min(c_{1},c_{2})$. DiTASK adapts ${\mathbf{W}}$ by learning transformations on the singular values $\mathbf{\Sigma}$ using a small set of parameters ${\bm{\theta}}$. This reduces the gradient storage requirement to ${\mathcal{N}}_{\mathcal{P}}+p+c_{1}c_{2}$, where ${\mathcal{N}}_{\mathcal{P}}$ is the number of intervals over which the CPA velocity field is defined. Unlike LoRA, DiTASK avoids gradients tied to low-rank matrices, significantly reducing memory usage for tasks with high-rank requirements or large input dimensions.
@@ Appendix D Additional Ablations
VTAB Benchmark. To understand the single-task generalization capabilities of our method, we compare DiTASK and LoRA [19] using the ViT [12] backbone on the VTAB Benchmark [59]. Shown in Figure 9, DiTASK achieves competitive performance with an order of magnitude fewer learnable parameters, presenting itself as a strong fine-tuning method for vision transformers.
@@
Additional Experiments. To understand the robustness of our design for different backbones, we evaluate DiTASK using the Pyramid Vision Transformer (PVT) [51] backbone against various parameter budgets of MTLoRA and LoRA. From Table 6, we conclude that our DiTASKachieves 2.5$\times$ improvement in multi-task performance over best-performing baseline using 4.4$\times$ fewer trainable backbone parameters.